### v1

# Vision-driven Automated Mobile GUI Testing via Multimodal Large Language Model

Abstract
With the advancement of software rendering techniques, GUI pages in mobile apps now encompass a wealth of visual information, where the visual semantics of each page contribute to the overall app logic, presenting new challenges to software testing. Despite the progress in automated Graphical User Interface (GUI) testing, the absence of testing oracles has constrained its efficacy to identify only crash bugs with evident abnormal signals. Nonetheless, there are still a considerable number of non-crash bugs, ranging from unexpected behaviors to misalignments, often evading detection by existing techniques. While these bugs can exhibit visual cues that serve as potential testing oracles, they often entail a sequence of screenshots, and detecting them necessitates an understanding of the operational logic among GUI page transitions, which is challenging traditional techniques. Considering the remarkable performance of Multimodal Large Language Models (MLLM) in visual and language understanding, this paper proposes a vision-driven automated GUI testing approach VisionDroid to detect non-crash functional bugs with MLLM. It begins by extracting GUI text information and aligning it with screenshots to form a vision prompt, enabling MLLM to understand GUI context. The function-aware explorer then employs MLLM for deeper and function-oriented GUI page exploration, while the logic-aware bug detector segments the entire exploration history into logically cohesive parts and prompts the MLLM for bug detection. We evaluate VisionDroid on three datasets and compare it with 10 baselines, demonstrating its excellent performance. The ablation study further proves the contribution of each module. Moreover, VisionDroid identifies 29 new bugs on Google Play, of which 19 have been confirmed and fixed.
Index Terms:
Automated GUI testing, Android app
I Introduction
In an era where mobile applications are integral to everyday life, ensuring the quality and reliability of these applications is paramount [1]. The field of mobile app development is characterized by rapid iterations and the need for continuous integration and deployment, making the quality assurance process both critical and challenging. As one kind of acceptance testing, automated Graphical User Interface (GUI) testing [2, 3, 4, 5, 6, 7, 8, 9, 10] has become a cornerstone in this process, promising to streamline testing by systematically identifying issues that could degrade user experience.
Despite the advancements in automated testing techniques, the intricate nature of mobile interfaces, coupled with the variability of user interactions, often necessitates the human testers as the last line of defense in ensuring app quality. Besides severe app crashes with obvious oracle signals, many bugs present visual cues such as misalignments, unintended element overlaps, or deviations from expected results which current automated systems frequently overlook [11]. As demonstrated in Fig. 1, the first sub-figure exposes a bug where the total income (5,030) does not match the summarization of the detailed records. Additionally, the four sub-figures together unveil a bug related to the failure of adding a record, as the income-expense overview in the fourth sub-figure should display five records after adding a new one. Our empirical study in Section II identifies nearly nine categories of such bugs, indicating the diversity of these non-crash bugs with visual cues. Although these bugs exhibit visual cues that make visual detection possible, they often entail a sequence of screenshots (as demonstrated in Fig. 1), requiring the understanding of the overall operational logic among these GUI page transitions. Only human testers with both sharp eyes and knowledge from extensive training can identify those bugs.
Figure 1: Example of bugs with visual cues
However, human testing is time-consuming and suffers from inherent scalability bugs. Although researchers begin to explore detecting non-crash or functional bugs with visual information, most of them are typically designed to identify specific types of bugs, such as display issues [12, 13], scaling problems [14, 15], game GUI glitches [16] or WebView bugs [17]. What makes it worse is that all of them require either a set of rules that are easy to be out of date, or a large amount of training dataset which takes much human effort for annotation. This specialization leaves a gap in the testing landscape for a more generalized, lightweight and intelligent approach that can adapt to the diverse challenges of mobile app testing especially for non-crash bugs.
Thanks to the emerging Multimodal Large Language Model (MLLM) [18], it is trained on the ultra-large-scale corpus and potentially able to understand the GUI pages as well as the information changes along with the page transition. Inspired by this, we propose a vision-driven automated GUI testing approach with MLLM. Nevertheless, three challenges need to be addressed. First, solely feeding the raw screenshot would be far less effective in letting the MLLM understand the GUI context, therefore how to align other sources of information and produce an effective vision prompt. Second, the triggering of bugs largely depends on whether and how the corresponding GUI page can be reached, therefore how to guide the MLLM to plan the exploration to delve deeper, covering diverse scenarios, and focusing on the app’s functionalities. Third, detecting the functional bugs requires tremendous domain knowledge and MLLM might lack it, therefore how to provide accurate information to facilitate MLLM making decisions.
To tackle these challenges, we propose VisionDroid, a vison-driven approach for detecting non-crash functional bugs with MLLM. It begins by extracting GUI text information and aligning it with GUI screenshots to form a vision prompt, enabling MLLM to understand GUI context. Second, the function-aware explorer utilizes the MLLM for GUI page exploration. It maintains a testing history comprising both the higher-level catalogue of testing progress and a detailed snapshot of each performed action, to enhance the exploration to be deeper and function-oriented. Third, the logic-aware bug detector breaks the entire exploration sequence into logically cohesive parts that mirror the app’s logical operational sequences. Based on them, the MLLM detects the bugs potentially by checking whether the GUI information changes between page transitions align with the processing logic. To boost the performance, we leverage in-context learning schema to provide useful examples when querying the MLLM, from online bug reports and history running records.
To evaluate the effectiveness of VisionDroid, we experiment on three datasets. Compared with 10 common-used and state-of-the-art baselines, VisionDroid can achieve more than 8%-230% and 27%-43% boost in average recall and precision compared with the best baseline, resulting in 50%-76% precision and 42%-64% recall. To further understand the role of each sub-module of the approach, we conduct ablation experiments to further demonstrate its effectiveness. We also evaluate the usefulness of VisionDroid by detecting unseen non-crash bugs in real-world apps from Google Play. Among 105 apps, VisionDroid detects 29 new crash bugs with 19 of them being confirmed and fixed by developers, while the remaining are still pending.
The contributions of this paper are as follows:
  • Vision. The pioneering work to formulate the non-crash functional bugs detection problem to an interactive multimodal question & answering task, enabling the MLLM to identify bugs by comprehending GUI visual information and functional logic.
  • Technique. A vision-driven automated GUI testing approach VisionDroid [^1] We release the source code, dataset, and detailed experimental results on our website https://github.com/testtestA6/VisionDroid. which designs a function-aware explorer and logic-aware bug detector for detecting non-crash functional bugs via MLLM.
  • Evaluation. Effectiveness and usefulness evaluation of the VisionDroid in the real-world apps with practical bugs detected (Section V-A and V-C).
II motivational study
To better understand the non-crash functional bugs in real-world practice, we carry out a pilot study to examine their prevalence. It also explores what categories of non-crash functional bugs, to facilitate the design of our approach.
II-A Data Collection
Our experimental dataset is collected from one of the largest open-source Android app hosting websites (GitHub) [19], which has a great number of Android bug reports. We employ the following criteria for app selection: more than 1K downloads in Google Play (popular), a public issue tracking system (traceable), and more than three years of development history (trustworthy). A total of 3,128 open-source Android apps with 57,326 bug reports are selected in this study. These apps belong to diversified categories such as business, communication, finance, etc. For the selected apps, we randomly select 6,000 bug reports involving 576 apps with detailed bug descriptions and GIFs/screenshots, and conduct manual analysis. Among them, 4,856 (81%) bug reports are related to the functional aspects of the app, while the rest are about the performance, security, etc. These functional bugs include 1,796 crash bugs (37%), and 3,060 non-crash functional bugs (63%). Since the crash bugs have been widely studied in previous work, we further investigate the non-crash functional bugs.
Figure 2: The categories of non-crash functional bugs
Figure 3: Overview of VisionDroid
II-B The Categories of Non-crash Functional Bugs
For the above-mentioned 3,060 non-crash functional bug reports, we manually classify them based on bug descriptions and reproduction steps, and derive the categories of non-crash functional bugs to attain a comprehensive understanding of them. Following the open coding protocol [20], two authors with more than three years of development and testing experience individually examine the issue description and reproduction steps. Then each annotator iteratively merges similar codes, and any disagreement of the categorization will be handed over to the third researcher for double-checking. Finally, we come out with a categorization of the non-crash functional bugs within (intra-page) and among the pages (inter-page), with details summarized in Fig. 2. We additionally provide examples for each category in our website1.
Intra-page non-crash functional bug. Intra-page non-crash functional bug denotes these bugs only related to a single GUI page, mainly involving the issues about information display as shown in Fig. 2.
Inter-page non-crash bug. Inter-page non-crash functional bug denotes these bugs related to multiple GUI pages, i.e., involving a testing sequence to trigger it, with details in Fig. 2. Among them, data operation failures mainly manifested as failures in adding, deleting, and modifying data, such as failed user information updates, errors in adding or removing shopping cart items, etc. Media control no-response mainly involves issues with video and audio playback, such as no response of playback control (failure of play, pause, stop functions). Numerical calculation errors mainly include mathematical errors, inaccurate statistical data, and cost calculation deviations. Setting and configuration failures mainly manifested as users being unable to save personalized settings or configurations, including ineffective or reset settings within the app (such as notification switches and theme changes). Navigation and link errors mainly manifested as page redirection errors and confusion in navigation logic between functions.
Findings. The above analysis confirms the popularity of non-crash bugs in mobile applications, implying the need for comprehensive testing. Meanwhile, considering the diversity of application functionality and bug manifestation, manually designing bug detection rules and building a complete set of defect methods require significant effort, and the automated technique is highly demanded. Despite their complexity, we have observed that these non-crash functional bugs typically manifest visual cues, either within a single GUI page or across a sequence of GUI pages. This motivates us to apply visual-driven techniques for bug detection.
III Approach
This paper proposes VisionDroid, a vision-driven automated GUI testing approach for detecting non-crash functional bugs with the Multimodal Large Language Model (MLLM). As shown in Fig. 3, it begins by extracting GUI text information, and aligning it with GUI screenshots, forming the vision prompt for GUI context perception. Second, the function-aware explorer utilizes the MLLM for guiding the exploration to delve deeper, encompassing more diverse pages, and focusing on the app’s functionalities. This is facilitated through maintaining the testing history comprising both the higher-level catalog of testing progress and the detailed snapshot of each performed action. Third, the logic-aware bug detector breaks the entire exploration sequence into locally cohesive parts that mirror the app’s logical operational sequences. Based on them, the MLLM detects the bugs potentially by checking whether the GUI information changes between page transitions align with the processing logic. To further boost the performance, we leverage the in-context learning schema to provide useful examples when querying the MLLM, from online issue reports and history running records.
III-A Alignment between Text and Image
Existing automated GUI testing approaches typically rely on view hierarchy files to obtain the attribute information (text, coordinates, actions, etc.) of app widgets to capture the current context and determine which widgets should be acted upon. This could be less efficient since the conversion from screenshot image to text could potentially lose information about the structure semantics of the GUI. Thanks to the MLLM, we can directly feed the screenshot image as input. Yet solely based on the screenshot also has its limitations since the image tends to utilize icons and extremely concise text for conveying information, which can also hinder the MLLM’s perceptions. Therefore, we design the visual and text alignment method, aiming at enhancing the MLLM’s understanding of the semantics of the GUI page.
III-A1 GUI Text Information Extraction
We extract three types of information, i.e., the app, the current GUI page, and all its contained widgets. The app information is extracted from the AndroidManifest.xml file, while the other two types of information are extracted from the view hierarchy file of the current GUI page, which can be obtained by UIAutomator [21]. We first extract the app information including the app name, and the activity names in this app , which can facilitate the MLLM to gain a general perspective of the functions. Then, we extract the page information from the view hierarchy file and extract the widgets information including “text” field, “hint-text” field, “resource-id” field, “class” field, “clickable” field, “long-clickable” field, “checkable” field, “scroll” field and “bounds” field, which facilitates the MLLM in suggesting actions related to these widgets.
Figure 4: Example of alignment text with image
III-A2 Text-Image Alignment for Visual Prompt
We annotate the GUI pages to guide the LLM’s attention towards the actionable parts (i.e., widgets). as shown in the right part of Fig. 4. Following existing studies [6, 7, 8, 9, 10], we support five types of basic operational actions, i.e., click, input, long-click, check, and scroll. We draw bounding boxes for each of these operational widgets, which are based on the extracted widget coordinates from the property of the “bounds” field, and use different colors indicating different types. In detail, if the “clickable” property of a widget is “true”, we draw a red bounding box to indicate it can be clicked, and if the “class” property is “EditText”, we draw a blue bounding box to indicate it needs input text. Similarly, we respectively use the purple, pink and yellow bounding boxes to indicate the widget with “long-clickable”, “checkable”, and “scrollable” properties as “true”. Inspired by the screen reader [22], we reconstruct the coordinate system in order from top to bottom and from left to right, and label each action widget on the image using Arabic numerals starting from 1. To mitigate the crowded appearance resulting from the abundance of components on a GUI page, we only mark the first widget of each row and let the MLLM infer others as shown in the prompt on the right side of Fig 4.
We then associate each annotated widget in the screenshot with the extracted GUI text information. Note that, in the “prompts related to GUI text and visual information” (details in Section III-B2 and Fig. 4), we would not provide the detailed alignment information considering the input length constraints of the MLLM. Instead, this alignment information will be used in “prompts related to action & feedback” (details in Section III-B2 and Fig. 5) to help verify whether the MLLM is suggesting interactive widgets in accordance with the correct logical inference process, akin to an explicit chain-of-thought mechanism. In detail, when querying the general action or text input, we not only request the MLLM to provide the widget ID from the screenshot, but we also ask for the corresponding widget text. When these two sources of information do not align, we infer potential errors in the MLLM’s decision-making process and prompt it to re-try.
III-B Function-aware Explorer
The non-crash functional bugs are often linked to the functionalities of the app, which often require lengthy operational sequences to trigger. Consequently, we design the function-aware explorer to guide the exploration of the app’s functionalities, while also encouraging deeper dives and broader coverage of GUI pages.
Figure 5: Example of function-based history
III-B1 Function-based Testing History Construction
To facilitate exploration of the app’s functionality, it is important to know the testing process, as the functions are usually accomplished by following a sequence of GUI pages. We maintain a complete testing history database, including textual descriptions and screenshots, and offer two views supporting the follow-up exploration and bug detection, as demonstrated in Fig. 5.
The TEXT view depicts the higher-level catalog of testing progress, organized according to the functions. It includes a list of tested functions along with their visit counts, which can assist the MLLM in avoiding duplicate exploration and uncovering untouched functions. To accomplish this, during the testing process, the MLLM is prompted to infer what function is currently being tested through the “prompt of function inquiry”, as demonstrated in Fig 5 (b). This enables us to gather statistics on the tested functions
The IMAGE view offers a detailed snapshot of the testing records through screenshots, providing a granular view of the testing process. Specifically, it includes a list of detailed GUI pages, with each page marked with a widget (or a group of widgets), i.e., denoting the corresponding actions when exploring that GUI page, which is demonstrated in Fig. 5 (a). Additionally, we provide derived functions for each screenshot to further ensure alignment between visual and textual information. We also associate a unique identifier (ID) with the function name, representing different pages/steps within the same function.
Figure 6: Example of explorer prompt
III-B2 Explorer Prompt Generation
With the static GUI information (in Section III-A1) and the dynamic testing progress (in Section III-B1), we devise prompts for inputting into the MLLM.
Prompts related to GUI text and visual information are demonstrated in Fig. 6 with “GUI text info” and “legend of current GUI screenshots”, offering the MLLM an overview of the current GUI page.
Prompts related to action & feedback are illustrated in Fig. 6 with “querying general action”, “querying text input”, and “testing feedback”. We prompt the MLLM to identify what operation is required, and provide the output template to assist the MLLM in generating a desired executive command for app testing. It’s worth noting that we separate the patterns for querying general action (e.g., click) and text input (e.g., input certain text) respectively, to enable the generated commands to better match the widgets in the GUI page. For the feedback question, if we detect the widget ID does not correspond to its textual description, which is determined with the help of the alignment information in Section III-A, we inform the MLLM that there might be an error in its decision-making process and prompt it to retry. ‘widget number’ is not accurate, better to use ‘widget id’, and revise them in other places and the figures.
Prompts related to function-based history are depicted in Fig. 6 with “explored functions” and “legend of recently tested screenshots”. They are fetched from the testing history, where the former is directly obtained from the text view, and the latter is extracted from the image view, displaying the latest four tested GUI screenshots to provide the most relevant clues.
Prompts related to function inquiry are illustrated in Fig. 6 with “querying function”, prompting the MLLM to identify what function is currently tested, and the prompt also provides an output template to assist the MLLM in generating the function name and its status.
Figure 7: Example of sequence segmentation and bugs detected by VisionDroid
III-C Logic-aware Bug Detector
During the exploration process, intra-page non-crash bugs are directly detected using MLLM, triggered by the “bug detect” prompt shown in Fig. 6. This section will focus on describing the bug detector for inter-page non-crash bugs.
Once the exploration phase is complete, we leverage the entire exploration history for bug detection. However, the overall history is often too lengthy, intricate, and noisy for directly querying the MLLM. The direct manner is to organize the entire history into individual parts based on the functions inferred by MLLM during the exploration process. Yet the MLLM tends to yield relatively granular function names, often comprising only 2 to 3 operational steps. However, many functional bugs are typically associated with the business logic formed by such consecutive functions. For instance, as depicted in Fig. 7, a bug involves functions like “edit expense”, “delete a budget”, “check budget”, etc. To tackle this, we propose a logic-driven testing sequence segmentation method to partition the entire exploration history into logically cohesive parts.
III-C1 Logic-driven Testing Sequence Segmentation Algorithm
The principle of test sequence segmentation is that GUI pages with similar operational logic are considered as a sub-sequence, while GUI pages with different logic are considered as different sub-sequences. Based on this, we begin by constructing a graph based on the entire testing history. Each node represents an explored GUI page (i.e., a screenshot from the image view in the testing history as shown in Fig. 5). Each edge signifies the transition between nodes. We assign the edge weights based on the semantic similarity of their associated function names (i.e., the text below each screenshot as shown in Fig. 5 (b)). Specifically, we segment the textual function name into tokens, and calculate the cosine similarity between them. In this way, the edge weight between pairs of pages with completely different functions should be 0, e.g., Check budget-2 and Setting-1 in Fig. 7.
Based on the weighted directed graph, the logic-driven testing sequence segmentation algorithm is conducted with the Louvain algorithm [23] and optimized based on modularity $Q$ for sub-sequence partitioning.
$Q=\frac{1}{2m}\sum_{ij}\left[A_{ij}-\frac{1}{2m}\right]\delta(c_{i},c_{j})$
$A_{ij}$ represents the weight between node $i$ and node $j$. $m$ is the sum of the weights of the graph. $c_{i}$ and $c_{j}$ are the sub-sequence number for node $i$ and node $j$. $\delta$ is an indicator function, when $c_{i}=c_{j}$ the value is 1, otherwise, it is 0, used to determine whether two nodes belong to the same sub-sequence.
Figure 8: Example of bug detection prompt
III-C2 Detector Prompt Generation
With the segmented results, we proceed to feed each sub-sequence into the MLLM for bug detection. Specifically, we provide MLLM with the related screenshots in the form of IMAGE view as shown in Fig. 5.
Prompts related to testing sub-sequence are demonstrated in Fig. 8 with “APP info” and “legend of GUI screenshots with function name”, offering the MLLM with app’s overview and the details about the derived testing sub-sequence.
Prompts related to bug detection question are demonstrated in Fig. 8. We enable the MLLM to comprehend the semantics of the test sequence based on the provided information and assess the correctness of the test path, further inquiring if there are any bugs in the test sequence and identifying which step contains bugs.
III-D Enriching Detector Prompt with Examples
It is usually difficult for MLLM to perform as well on domain-specific tasks as ours, and common practice would be employing the in-context learning schema to boost the performance. It provides the MLLM with examples to demonstrate what the instruction is, which enables the MLLM to better understand the task. Following the schema, along with the prompt for the detector, as described in Section III-C, we additionally provide the MLLM with examples of the non-crash logical bugs.
To achieve this, we first build a basic example dataset of non-crash bugs from the issue reports of open-source mobile apps and continuously enlarge it. For each data instance, it records the bug description, bug screenshot and natural language described bug reproduction path which facilitates the MLLM understanding of what the non-crash bug is. We select 200 bugs (100 intra-page, 100 inter-page) with the highest quality and complete information as the example data. We also enrich the example dataset with the newly emerged non-crash bugs which truly trigger bugs during VisionDroid runs on various apps.
Although examples can provide intuitive guidance to the LLM in accomplishing a task, excessive examples might mislead the LLM and cause the performance to decline. Therefore, we design a retrieval-based example selection method to choose the most suitable examples (i.e., most similar to the tested GUI page) for MLLM. The similarity comparison is based on the activity name of the tested page. Similar to previous works [24], We use Word2Vec [25] to encode the context information of each activity into a sentence embedding, and calculate the cosine similarity between the activity and each data instance in the example dataset. We choose the top-K data instance with the highest similarity score, and set K as 5.
III-E Implementation
We implement VisionDroid based on the gpt-4 API which is released on the OpenAI website [^2] https://beta.openai.com/docs/models/GPT-4. It obtains the view hierarchy file of the current GUI page through UIAutomator [21] to extract text information of the input widgets. VisionDroid is implemented as a fully automated GUI testing tool, which uses or extends the following tools: VirtualBox and the Python pyvbox [26] for running and Android Debug Bridge (ADB) [27] for interacting with the app under test.
IV Experiment Design
  • RQ1: (Coverage and bugs detection performance) How effective of VisionDroid in testing coverage and bug detection?
For RQ1, we first present some general views of VisionDroid for coverage of the explorer and bug detection of the detector. Then we compare VisionDroid with commonly-used and state-of-the-art baseline approaches.
  • RQ2: (Ablation Study) What is the contribution of the (sub-) modules of VisionDroid for coverage and bug detection performance?
For RQ2, We conduct ablation experiments to evaluate the impact of each (sub-) module on the performance.
  • RQ3: (Usefulness Evaluation) How does our proposed VisionDroid work in real-world situations?
For RQ3, we use VisionDroid to automatically explore the app, and detect unseen non-crash functional bugs, and issue the detected bugs to the development team.
IV-A Datasets
For RQ1-RQ2, the experimental dataset comes from three sources. The first is from the Odin and RegDroid datasets [28, 29] (We name them baseline dataset), which contains 131 non-crash functional bugs that can be reproduced in our experimental environment.
Considering the small number of apps in the baseline dataset, we collect a second dataset (Github dataset) following similar procedures as the baseline dataset. In detail, we crawl the 30 most popular apps of each category from Google Play [30], and we keep the ones with at least one update after May. 2024, resulting in 218 apps in 10 Google Play categories. Then we manually review each issue report and app, and filter them according to the following criteria: (1) the app wouldn’t constantly crash on the emulator; (2) it can run all baselines; (3) UIAutomator [21] can obtain the view hierarchy file and its GUI screenshot; (4) the issue report involve non-crash bug and it can be manually reproduced for validation; (6) the app is not used in the motivational study or example dataset construction. Please note that we follow the name of the app to ensure that there is no overlap between the datasets. Finally, 43 apps with 247 non-crash bugs remain for further experiments.
To avoid potential data leakage issues caused by the use of MLLM and further analyze the bug detection capability of VisionDroid, we construct a third dataset (Injection dataset) through bug injection, which can ensure that MLLM has not learned about these bugs in the open-source repositories. Specifically, we obtain the source code of the second dataset apps from GitHub and manually verify whether the app could conduce the bug injection by modifying the code through repackaging. Among them, 27 apps can run normally after bug injection. We conduct the injection based on the bug description, bug repair code, and bug root cause obtained through the pilot study in Section II, and implement it following the existing bug injection approach [31, 32, 33]. For the 9 summarized bug categories in Section II, we inject 5 non-crash bugs per category, resulting in a total of 45 bugs in 27 apps.
TABLE I: Result of coverage and bug detection (RQ1)
Method Baseline data GitHub data Bug injection data
Act Cod Rep Bug P R Act Cod Rep Bug P R Act Cod Rep Bug P R
Intra-page bug detection method with automated GUI testing
OwlEyes+A 0.29 0.27 92 18 0.20 0.14 0.41 0.37 117 57 0.49 0.23 0.42 0.40 77 7 0.09 0.16
NightHawk+A 0.29 0.27 86 21 0.24 0.16 0.41 0.37 99 66 0.67 0.27 0.42 0.40 41 8 0.20 0.18
DiffDroid+A 0.29 0.27 87 20 0.23 0.15 0.41 0.37 82 47 0.57 0.19 0.42 0.40 27 3 0.11 0.07
dVermin+A 0.29 0.27 67 6 0.09 0.05 0.41 0.37 51 10 0.20 0.04 0.42 0.40 18 2 0.11 0.04
OwlEyes+F 0.39 0.37 107 17 0.16 0.13 0.44 0.40 119 57 0.48 0.23 0.43 0.39 79 8 0.10 0.18
NightHawk+F 0.39 0.37 100 20 0.20 0.15 0.44 0.40 98 64 0.65 0.26 0.43 0.39 39 9 0.23 0.20
DiffDroid+F 0.39 0.37 94 20 0.21 0.15 0.44 0.40 77 51 0.66 0.21 0.43 0.39 28 4 0.14 0.09
dVermin+F 0.39 0.37 41 8 0.20 0.06 0.44 0.40 44 13 0.30 0.05 0.43 0.39 11 2 0.18 0.04
OwlEyes+H 0.38 0.35 97 12 0.12 0.09 0.44 0.40 97 54 0.56 0.22 0.43 0.39 71 7 0.10 0.16
NightHawk+H 0.38 0.35 88 16 0.18 0.12 0.44 0.40 94 58 0.62 0.23 0.43 0.39 69 7 0.10 0.16
DiffDroid+H 0.38 0.35 75 17 0.23 0.13 0.44 0.40 78 32 0.41 0.13 0.43 0.39 31 4 0.13 0.09
dVermin+H 0.38 0.35 33 5 0.15 0.04 0.44 0.40 27 17 0.63 0.07 0.43 0.39 11 2 0.18 0.04
OwlEyes+G 0.47 0.43 99 17 0.17 0.13 0.49 0.44 104 59 0.57 0.24 0.47 0.41 77 8 0.10 0.18
NightHawk+G 0.47 0.43 95 18 0.19 0.14 0.49 0.44 110 69 0.63 0.28 0.47 0.41 48 9 0.19 0.20
DiffDroid+G 0.47 0.43 85 16 0.19 0.12 0.49 0.44 88 49 0.56 0.20 0.47 0.41 27 3 0.11 0.07
dVermin+G 0.47 0.43 37 6 0.16 0.05 0.49 0.44 29 18 0.62 0.07 0.47 0.41 13 3 0.23 0.07
Intra-page bug detection method
iFixdataloss 0.33 0.30 21 8 0.38 0.06 0.44 0.40 52 18 0.35 0.07 0.43 0.39 13 2 0.15 0.04
SetDroid 0.31 0.27 22 8 0.36 0.06 0.44 0.40 98 44 0.45 0.18 0.43 0.39 11 2 0.18 0.04
Genie 0.32 0.31 55 12 0.22 0.09 0.44 0.40 61 22 0.36 0.09 0.43 0.39 27 4 0.15 0.09
Odin 0.33 0.30 57 22 0.39 0.17 0.44 0.40 58 29 0.50 0.12 0.43 0.39 29 6 0.21 0.13
Lint - - 1 1 1.00 0.01 - - 2 1 0.50 0.00 - - 1 1 1.00 0.02
LiveDroid - - 1 1 1.00 0.01 - - 3 1 0.33 0.00 - - 2 0 0.00 0.00
GPT-4 0.39 0.31 55 12 0.22 0.09 0.44 0.40 61 22 0.36 0.09 0.43 0.39 18 2 0.11 0.04
VisionDroid 0.57 0.55 109 55 0.50 0.42 0.65 0.56 177 127 0.72 0.51 0.61 0.59 38 29 0.76 0.64
  • Notes: “Act” is activity coverage, “Cod” is Code coverage. “Rep” is number of bugs reported by tools, “Bug” is true bug in the reported bug, “P” is precision and “R” is the recall. “+A” means it with APE, “+F” means it with Fastbot, “+H” means it with Hunmanoid, and “+G” means it with GPTDroid.
For RQ3, we further evaluate the usefulness of VisionDroid in detecting unseen bugs. We randomly selected 105 popular apps based on app type and download number. We use the same configurations as the previous experiments. Once a bug is spotted, we create an issue report by describing the bug, and report them to the app development team through the issue reporting system or email.
IV-B Baselines
To demonstrate the advantage of VisionDroid, we compare it with 10 common-used and state-of-the-art non-crash bug detection techniques and 4 common-used automated GUI testing techniques. We classify these baselines based on the categories of bugs, including intra-page bug detection approaches and inter-page bug detection approaches.
For intra-page non-crash bug detection, the techniques are mainly for GUI display-related issue detection. OwlEyes [12] and NightHawk [13] use deep learning to detect UI display issues (e.g., overlapping texts, component occlusion). DVermin [15] detects scaling issues by comparing GUI pages under different page settings. DiffDroid [14] compares the GUI pages of an app on two different devices to find compatibility issues. The input of these techniques are screenshots/visual hierarchy files and need to be integrated with automated testing approaches to fetch the required information during the exploration. We utilize four popular automated GUI testing tools, i.e., model-based tools APE [9] and Fastbot [34], learning-based tool Humanoid [35] and LLM-based GPTDroid [24].
For inter-page non-crash bug detection, we choose 6 baselines from different perspectives. iFixdataloss [36] is the most recent work of finding data loss issues. SetDroid [37] uses metamorphic testing to find system setting-related functional bugs. Genie [38] uses the metamorphic relation of independent view property to find non-crash functional bugs, while Odin [29] uses the implicit knowledge of “bugs as deviant behavior” to find non-crash functional bugs. Android Lint [39] is a popular static analysis tool for finding different issues in Android apps based on predefined rules. LiveDroid [40] is also a static analysis tool for detecting data loss issues.
IV-C Experimental Setup
We deploy the baselines and our approach on a 64-bit Ubuntu 18.04 machine (64 cores, AMD CPU, 2GB RAM) and evaluate them on Google Android 7.1 emulators. Following common practice [7, 8, 10], we registered separate accounts for each bug that requires login and wrote the login scripts, and during testing reset the account data before each run to avoid possible interference. To ensure fair and reasonable use of resources, we set up the running time of each tool in one app to 50 minutes, which is widely used in other GUI testing studies [8, 10, 41, 34] We run each tool 3 times and take the average value as the result to mitigate potential bias [8, 10, 41, 34].
TABLE II: Contribution of different sub-modules (RQ2)
Sub-modules Baseline data GitHub data Bug injection data
Act Cod Rep Bug P R Act Cod Rep Bug P R Act Cod Rep Bug P R
VisionDroid 0.57 0.55 109 55 0.50 0.42 0.65 0.56 177 127 0.72 0.51 0.61 0.59 38 29 0.76 0.64
w/o Alignment 0.44 0.39 97 32 0.33 0.24 0.49 0.48 133 75 0.56 0.30 0.47 0.43 29 18 0.62 0.40
w/o History 0.41 0.35 91 28 0.31 0.21 0.47 0.44 128 63 0.49 0.26 0.44 0.31 27 15 0.56 0.33
w/o Community 0.57 0.55 90 17 0.19 0.13 0.65 0.56 149 39 0.26 0.16 0.61 0.59 42 7 0.17 0.16
w/o Example 0.57 0.55 89 18 0.20 0.14 0.65 0.56 151 41 0.27 0.17 0.61 0.59 39 8 0.21 0.18
IV-D Evaluation Metrics
We measure the performance from coverage and bug detection, respectively corresponding to the explorer and detector in VisionDroid. For coverage, we obtain the activity coverage and code coverage [42, 43, 44, 10, 45, 35], in which treat the activities defined in the AndroidManifest.xml file of an Android app as the whole set of activities [46, 47, 7].
For bug detection, following existing studies [38, 29], we report the detected bugs by VisionDroid, true bugs among the detected bugs. Since we know the entire set of bugs, we also calculate the precision and recall of the bug detection.
V Results and Analysis
V-A Coverage and Bug Detection Performance (RQ1)
We present the coverage and bug detection performance of VisionDroid and baselines, in terms of the baseline dataset, GitHub dataset and the bug injection dataset in Table I.
We first present the coverage of VisionDroid, which can cover far more activities and code than the baselines. In three datasets, the average activity coverage is between 57% and 61%, and the average code coverage is between 55% and 59%. It achieves between 21% (0.57 vs. 0.47) to 33% (0.65 vs. 0.49) activity coverage and between 27% (0.56 vs. 0.44) and 43% (0.59 vs. 0.41) code coverage higher even compared with the best baseline (GPTDroid) across three datasets. This indicates the effectiveness of VisionDroid in covering more activities and codes, thus bringing higher confidence to app quality, potentially uncovering more bugs.
The average precision and recall of VisionDroid in three datasets achieve 50%-76%, and 42%-64%, respectively. It achieves between 8% (0.72 vs. 0.67) and 230% (0.76 vs. 0.23) precision and between 27% (0.56 vs. 0.44) and 43% (0.59 vs. 0.41) higher even compared with the best baseline across three datasets. Please note that although some baseline methods can achieve higher precision, they detect fewer bugs. In terms of the number of detected bugs, VisionDroid can still achieve a higher detection rate. We also compare the similarities and differences of the bugs between the best baselines (Odin and NightHawk) and our approach, and the results show that all bugs detected by them are also detected by VisionDroid. This indicates the effectiveness of VisionDroid in detecting bugs and helps to ensure app quality.
The advantages of VisionDroid are due to three aspects. First, VisionDroid designs the vision prompt which can better capture the GUI context, while the baselines mainly rely on the view hierarchy file in which the widgets might have meaningless “text” or “resource-id”, hindering them from effectively understanding the GUI pages. Second, VisionDroid plans the exploration path both considering the semantics of a single GUI page and the exploration history which can enable the function-aware exploration, while the baselines can hardly understand the inherent business logic of the app, resulting in aimless exploration. Third, VisionDroid detects the bugs by understanding the semantics of a sequence of GUI pages, while the baselines can only detect specific types of bugs or have low detection accuracy.
Fig. 7 demonstrates an example of VisionDroid’s detected bug. This bug is that after deleting a single expense on the budget information page, all original expense records disappear when viewing the expense page. We can see that VisionDroid is adept at determining whether the functional logic of the app is abnormal based on the semantics of test path and GUI information.
V-B Ablation Study (RQ2)
V-B1 Contribution of Sub-modules
Table II further demonstrates the performance of VisionDroid and its 4 variants. We remove each sub-module of the VisionDroid in Fig. 3 separately, i.e., Alignment text with image, Function-based testing history construction, Function-driven community detection and Example bugs retrieval. For removing the Alignment text with image, we directly input the unmarked screenshot. For removing function-based testing history construction, we delete the history information. For removing the community detection, we randomly segment the test sequence and for removing retrieved examples, we use the random selection.
The experiment results (Table II) demonstrate that removing any of the sub-modules would result in a noticeable performance decline, indicating the necessity and effectiveness of the designed sub-modules. Removing the community detection (w/o-Community) has the greatest impact on the performance, reducing the bug detection performance by 347% (0.17 vs. 0.76) precision and 300% (0.16 vs. 0.64).
We also notice that, when removing the alignment text with image (w/o-Alignment) and function-based testing history construction(w/o-History), the activity coverage has an impact, reducing the activity coverage by 39% (0.41 vs. 0.57) and 57% (0.35 vs. 0.55). This further indicates that the alignment text with image can help MLLM understand the structure and semantic information of GUI pages and make reasonable judgments. The history can help retain knowledge during testing and gain global viewpoints to reach uncovered areas.
Besides, randomly selecting examples (w/o-retriExample) would also largely influence the performance, and decrease the bug detection performance by 262% (0.21 vs. 0.76) precision and 256% (0.18 vs. 0.64) recall. This indicates that by providing similar examples, the MLLM can quickly think out what should the no-crash bugs look like, which further indicates the demonstration can facilitate the MLLM in producing the required output.
V-B2 Influence of Different Number of Examples
Fig. 9 demonstrates the performance under the different number of examples provided in the prompt. We can see that the non-crash bug detection performance increases with more examples, reaching the highest bug detection rate with 4 examples. And after that, the performance would gradually decrease even increasing the examples. This indicates that too few or too many examples would damage the performance, because of the tiny information or the noise in the provided examples.
Figure 9: Different number of examples (RQ2)
V-C Usefulness Evaluation (RQ3)
For the 105 apps, VisionDroid detects 83 non-crash bugs in 53 apps, of which 37 bugs were never discovered before. Furthermore, only 9 of these new bugs are detected by Genie [38] and Odin [29]. We submitted these 37 bugs to developers, and 19 of them have been fixed/confirmed so far (10 fixed and 9 confirmed), while the remaining are still pending (none of them is rejected). This further indicates the effectiveness of our proposed VisionDroid in bug detection. Due to space limit, Table III presents these fixed/confirmed bugs, and the full lists can be found on our website1. We further analyze the details of these bugs, and 21 of them involve intra-page bugs and 16 of them involve inter-page bugs.
TABLE III: Confirmed or fixed bugs. (RQ3)
Id APP Name Category Download Status
1 MOMPA Music 500M+ fixed
2 Calendar Tool 100M+ fixed
3 DigiCal Tool 100M+ fixed
4 Expensify Finance 100M+ fixed
5 TrackWallet Finance 10M+ confirmed
6 Roadie Travel 10M+ confirmed
7 ISAY Commun 10M+ fixed
8 EasyBudget Finance 10M+ confirmed
9 Iplan Travel 10M+ fixed
10 MediaFire Product 5M+ confirmed
11 SmartGuide Navig 500K+ fixed
12 MMDR Utilities 500K+ confirmed
13 Sygic Travel 500K+ fixed
14 Fair Health 500K+ confirmed
15 ClassySha Tool 500K+ fixed
16 Linphone Commun 50K+ confirmed
17 Rocket Finance 50K+ confirmed
18 Monefy Finance 50K+ fixed
19 Origin Finance 50K+ confirmed
VI Related Work
VI-A Automated GUI Testing
To ensure the quality of mobile apps, many researchers study the automatic generation of large-scale test scripts to test apps [48]. Since Android apps are event-based [49, 50, 51, 52], the most common automated testing methods are model-based automated GUI testing methods [2, 3, 4, 5, 6, 7, 8, 9, 10], design corresponding models through the analysis of the apps. Due to the lack of consideration for the semantic information of the app’s GUI pages, the coverage of model-based methods is still low. Researchers further proposed human-like testing strategies and designed learning-based automated GUI testing methods [35, 46]. GPTDroid [24] used GPT-3.5 to generate the testing script, aiming at proposing a more effective approach to generate human-like actions for thoroughly and more effectively testing the app. However, these methods require first converting the GUI visual information into text, which can cause significant visual information loss during the process. Moreover, a large number of app widgets and pages are generated through dynamic rendering, and the attribute information of image widgets will be lost when obtaining view hierarchy files. As a result, the coverage of automated testing is low, and a large number of widgets and operations are missed. This study employs the MLLM for the GUI testing, which is more effective through aligning the GUI screenshot with text information for the vision prompt.
VI-B Non-crash Functional Bug Detection
According to the sources of test oracle, current research can be divided into four types. The first type utilizes oracle provided by testers [53] to find non-crash bugs in different strategies (such as ChimpCheck [54], AppFlow [55], ACAT [56], AppTestMigrator [57], CraftDroid [58]). Unlike them, this paper doesn’t require manual direct/indirect oracle.
The second type uses differential testing to overcome the problem of missing oracles. Genie [38] and Odin [29] used the metamorphic relation of independent view property to find bugs. SPAG-C [59], DiffDroid [14] and dVermin [15] compare screenshots between two different mobile devices or platform versions to identify bugs.
The third type is the automatic generation of test oracle for specific categories of user interactions. ITDroid [60] compared the GUI pages of an app under two different language settings to find i18n issues. SetDroid [37] generated the oracle for system settings bugs, but can’t detect other types of bugs.
The fourth type is to use labeled large-scale bug data to train computer vision models for detecting bugs related to GUI display. The OwlEyes [12], Gilb [16] and NightHawk [13] are based on large-scale labeled UI bug data, to train computer vision models that can be used to detect the display bugs.
However, these bug detection methods have problems such as a single category of bug detection, requiring a large amount of manually labeled test oracle data, and being unable to meet complex business scenarios and app functionality.
VII Conclusion
With the advancement of LLM, there have been tremendous applications of LLMs in software engineering-related tasks such as program repair and test case generation with great performance improvement. Subsequently, MLLMs emerge and are applied in image-text reasoning, video generation, etc., further expanding our understanding of the capabilities of LLMs. This paper proposes VisionDroid, the pioneering work using MLLMs for vision-based mobile app GUI testing, and the experiment demonstrates its effectiveness and usefulness. This work contributes as a stepping stone for future exploration in software engineering, offering insights and guidance on the potential applications of MLLM. In the future, we will further improve this approach for better performance. Meanwhile, beyond Android apps, we will customize and apply our approach to other platforms such as iOS, web, desktop and even embedded systems for software testing.